# astroCAMP: A Community Benchmark and Co-Design Framework for Sustainable SKA-Scale Radio Imaging

### v1

@@ 6.3. Weak CPU-only Scaling of WSClean+IDG
[…]
Relative to ideal scaling, CPU performance at 64 threads deviates by more than 95%. These results confirm that current software stacks do not scale linearly for SKA-sized workloads on modern CPU servers, underscoring the need for co-designed kernels and dataflows that are explicitly optimised for hierarchical memory systems and modern node/network topologies.
@@ 6.4. Location-Dependent Efficiency for SDPs
Figure 9. Carbon (left) and cost breakdown (right) with efficiency metrics for the largest image size (32768²) across all timesteps and channels configurations, comparing SKA-Low (WA) and SKA-Mid (SA).
In Fig. 9, we analyse the carbon and cost breakdown for our largest image size and relate it to performance efficiency at the two SKA SDP locations using site-specific grid carbon intensities and electricity prices. While throughput is identical at both sites, carbon and cost efficiencies diverge due to local grid and pricing conditions. In both locations, operational emissions dominate total carbon (90-95%), reflecting identical hardware and workload characteristics. Despite this, WA achieves higher carbon efficiency (Mvis/kgCO2) owing to its lower grid carbon intensity (0.321 vs. 0.672 kg CO2/kWh), indicating that throughput and workload characteristics, rather than grid intensity alone, primarily govern carbon efficiency.
In contrast, cost efficiency is primarily driven by CAPEX (89-92%). WA’s higher electricity cost (0.27 kWh vs. 0.19 kWh in WA) results in slightly lower cost efficiency (Mvis/\$) compared to SA, particularly for larger, longer-running workloads. Across both locations, efficiency improves with increasing timesteps and channels as embodied carbon and capital costs are amortized over greater computational output. This highlights that consolidating workloads into fewer, larger batch jobs can deliver meaningful sustainability and economic benefits, especially in regions with higher electricity prices or carbon intensities.
@@ 6.5. Design Space Exploration with PREESM
PREESM (PREESM2017) is a prime example of a DSE framework to study its extendability. The diversity of metrics, specialized for a scientific field, calls for user made metrics support. The current flow supports many system and hardware platform-level metrics, and more can be computed from those metrics based on user needs, using an existing model for fast evaluation.
Taking a step back to look at metric implementations, algorithmic metrics require the complete execution of the pipeline, with potentially long execution, incompatible with DSE. Future flow will require knowledge of parameters influence on expensive metrics to avoid systematic evaluation (honorat:hal-03752645). For example, a full execution of a pipeline to evaluate an algorithmic metric is needed when changing quantization parameters, but is not required when changing hardware frequency parameters.
@@
Sustainability and Economic metrics are two examples of environmental metrics, based on external information. Those metrics require interaction with external models or databases, bringing a potentially multi-model co-design flow.
@@ 7. Discussion of Benchmark Evaluation
The results presented in the previous section confirm the central hypothesis of this work: that only through algorithm–architecture co-design can radio interferometric imaging achieve sustainable exascale performance. Our preliminary evaluation of WSClean leads to three main conclusions:
@@
    Scalability limits: Current imaging algorithms do not exhibit ideal scaling at SKA problem sizes due to memory-bound operations and communication overheads.
@@
    Co-design imperative: Achieving sustainable scaling demands joint optimization of algorithms, hardware, and scheduling—targeting reduced-precision arithmetic, data locality, and carbon-aware workload management.
@@ 7.1. Why We Need HW–SW Co-Design?
Optimizing the SKA SDP requires the joint consideration of SW and HW design choices, as neither dimension alone is sufficient to fully exploit the available design space across relevant metrics such as energy, carbon emissions, cost, and execution time.
@@
This analysis demonstrates that SW and HW optimizations are inherently interdependent. SW-only approaches are constrained by inefficient HW utilization, while HW-only optimizations that target energy efficiency may have limited impact when dynamic energy is a small fraction of total energy. Effective optimization of sustainability metrics therefore requires a holistic HW–SW co-design approach, in which both dimensions are co-optimized to unlock their full potential.
7.2. Community Challenge: How Much Quality can we Trade off for Efficiency?
SKA documents define high-level goals—e.g. 1 % astrometry, spectral dynamic ranges of $10^{5}$$10^{4}$—but not survey-level tolerances on flux accuracy, completeness, PSF residuals, polarisation purity, or spectral-line fidelity. Without such thresholds, the co-design community cannot determine how much approximation (e.g., reduced precision, coarse $w$-stacking) is scientifically acceptable.
Open Challenge for the SKA Community For each SKA Key Science Programme—Cosmic Dawn/EoR, Galaxy Evolution and Cosmology, Cosmic Magnetism, and Time-Domain Astrophysics—define quantitative application-level quality metrics and tolerances. Examples include PSNR/SSIM, flux-scale accuracy, astrometric precision, polarisation purity, RM recovery error, spectral-line fidelity, time-series accuracy, and transient completeness. Tolerances must be precise enough to guide hardware–software co-design (precision choices, approximations, accelerators) while ensuring validity for: 21 cm EoR power spectra and tomographic cubes, all-sky continuum and deep extragalactic surveys, RM Grid measurements of cosmic magnetism, high-cadence transient and pulsar timing searches.
As SKA power and cost envelopes are fixed years in advance, the central question becomes: Which combinations of algorithmic parameters and hardware choices maximise performance and energy efficiency without violating scientific fidelity? Answering this requires cross-layer metrics connecting runtime, energy, carbon, utilisation, and image quality. But the community still lacks programme-level tolerances for key cosmological surveys (bacon2020cosmology).
@@ 8. Conclusions
This paper introduced an open benchmarking and co-design framework for sustainable, reproducible, and energy-efficient radio interferometry imaging at exascale. By coupling representative datasets with unified performance, energy, and carbon metrics, the framework provides a common ground for evaluating algorithms and architectures across heterogeneous platforms. Preliminary results confirm that existing imaging codes do not scale efficiently to SKA-class workloads and that energy and carbon constraints must drive future design choices. The proposed methodology promotes transparent, carbon-aware design-space exploration and establishes benchmarks as shared instruments for scientific accountability. In advancing co-design as both a technical and ethical imperative, this work contributes to building trust in high-performance computing and to reducing the environmental footprint of data-intensive science. We conclude with a community call to: (1) establish science-driven imaging-quality tolerances, (2) expand benchmark datasets, (3) develop reference submissions for current and emerging imaging pipelines, and (4) build an open, MLPerf-style (reddi2020mlperf; farrell2021mlperf) submission and validation process for the SKA era.

### v2

@@ 6.3. Strong CPU-only Scaling Limitations
Figure 13. Strong-scaling speedup and parallel efficiency for CPU-only WSClean vs GPU accelerated gridding.
Table 3. Execution times for $t=256$, $c=256$, $16384^{2}$ with IDG on CPU or GPU across CPU core counts.
Solution 1c 16c 32c 64c
WSClean + IDG CPU 3:13:10 1:08:28 1:08:10 1:08:06
WSClean + IDG GPU 1:09:27 1:09:57 1:09:14 1:08:45
Figure 14. Darshan summary for a CPU-only run (16 cores), showing I/O runtime share and operation counts.
Table 3 confirms this behavior on the flagship case: CPU-only runtime improves from one to 16 cores, then nearly plateaus, while GPU-enabled wall time remains nearly constant across CPU core counts because the dominant parallel work is already offloaded.
Although bandwidth-dominated kernels can strong-scale in principle; here, losses are dominated by serial regions and orchestration overheads in the full application path. Fig. 13 quantifies this collapse via speedup and parallel efficiency and indicates substantial synchronization overhead from four or more cores. Flamegraph traces also expose libcasa_tables and libcasa_casa in effectively single-threaded WSClean sections, creating a serial metadata/I-O lane that increases waiting time and weakens scaling gains from extra CPU cores. In our runs we use MSv2; moving to MSv4-capable I/O frameworks should reduce this serialization pressure (e.g., better concurrent access and request aggregation), improving pipeline-level scaling in addition to raw I/O efficiency.
Investigating further, we identify that the spin and synchronization overhead grows quickly with thread count, consistent with the observed speedup collapse. Fig. 14 shows that I/O accounts for a small fraction of total wall time ($\sim 1\%$), but with many short and frequently misaligned requests. For the largest case, we observe 99.66% misaligned file requests and 280,024 read operations below 1 MB (262,144 targeting the same data file), indicating that request aggregation and alignment remain relevant even when peak filesystem bandwidth is not saturated.
@@ 6.4. Location-Dependent Efficiency for SDPs
Figure 15. Carbon (top) and cost breakdown (bottom) with efficiency metrics for image size (32768²) across all configurations, comparing SKA-Low (WA) and SKA-Mid (SA).
In Fig. 15, we compare carbon and cost for the largest image size at the SKA SDP locations. Throughput is identical, but efficiency diverges with local grid intensity and electricity price. Operational emissions dominate carbon (90-95%); WA attains higher carbon efficiency (Mvis/kgCO2) because grid intensity is lower (0.321 vs. 0.672 kg CO2/kWh). Cost efficiency is CAPEX-dominated (89-92%): WA’s higher electricity price (0.27 vs. 0.19 \$/kWh in SA) yields lower Mvis/\$. Increasing timesteps/channels amortizes embodied carbon and CAPEX, so larger batches improve efficiency.
@@ 6.5. Design Space Exploration with PREESM
PREESM (Suriano et al., 2017) is a prime example of a DSE framework to study its extendability. The diversity of metrics, specialized for a scientific field, calls for automated exploration support. The current flow supports many system and hardware platform-level metrics, and more can be computed from those metrics based on user needs, using an existing model for fast evaluation.
Taking a step back to look at metric implementations, algorithmic metrics require the complete execution of the pipeline, with potentially long execution, incompatible with DSE. Future flows should model parameter influence on expensive metrics to avoid systematic evaluation (Honorat et al., 2022). For example, changing quantization may require full execution, while changing hardware frequency may not.
Sustainability and Economic metrics are two examples of environmental metrics, based on external information. These metrics require interaction with external models or databases, enabling a multi-model co-design flow.
@@ 7. Discussion of Benchmark Evaluation
The benchmark results support a clear result-to-decision chain for SKA imaging co-design. Our evaluation of WSClean+IDG yields two evidence-based conclusions:
@@
    Pipeline-level scaling limits dominate: CPU strong scaling collapses after low thread counts (Figs. 12, 13), CPU activity is bursty and mostly single-core in critical stages (Fig. 5), and end-to-end wall time remains far above summed kernel time even when active-window GPU utilization is nonzero (Fig. 9). Together with the roofline comparison (Fig. 10), this points to orchestration/synchronization and serial-path effects rather than a single kernel-bound limit.
@@
    Efficiency gains are configuration- and site-dependent: energy hierarchy and throughput trends (Figs. 6, 7) show accelerator-dominated scaling with persistent static-energy overheads; payload trends indicate input-dominated growth at large $(t,c)$ (Fig. 8); and carbon/cost efficiency diverges by location despite equal throughput (Fig. 15).
Practical takeaway: use astroCAMP in three steps: (i) detect scaling/utilization bottlenecks (Figs. 5, 12, 13, 9), (ii) compare energy/carbon/cost efficiency across $(t,c)$ and sites (Figs. 7, 15, 8), and (iii) select the smallest configuration that meets science-quality tolerances.
@@ 7.1. Why We Need HW–SW Co-Design?
7.2. Open Challenge for the SKA Community
How Much Quality can we Trade off for Efficiency? For each SKA Key Science Programme—Cosmic Dawn/EoR, Galaxy Evolution and Cosmology, Cosmic Magnetism, and Time-Domain Astrophysics—define quantitative application-level quality metrics and tolerances. Examples include PSNR/SSIM, flux-scale accuracy, astrometric precision, polarisation purity, RM recovery error, spectral-line fidelity, time-series accuracy, and transient completeness. Tolerances must be precise enough to guide hardware–software co-design (precision choices, approximations, accelerators) while ensuring validity for: 21 cm EoR power spectra and tomographic cubes, all-sky continuum and deep extragalactic surveys, RM Grid measurements of cosmic magnetism, high-cadence transient and pulsar timing searches.
Why is this challenge needed? SKA documents define high-level goals—e.g. 1 % astrometry and spectral dynamic ranges of $10^{5}$$10^{4}$—but not survey-level tolerances for flux accuracy, completeness, PSF residuals, polarisation purity, or spectral-line fidelity. Without these thresholds, the co-design community cannot determine which approximations (e.g., reduced precision, coarse $w$-stacking) are acceptable. With SKA power and cost envelopes fixed years ahead, the key question is which algorithm/hardware combinations maximize performance and energy efficiency without violating fidelity. This requires cross-layer metrics linking runtime, energy, carbon, utilization, and image quality, plus programme-level tolerances that are still missing (Bacon et al., 2020).
Community engagement plan. astroCAMP follows an open-source, maintainer-guided model, hosted as a public repository (SEAMS Project, 2025) where datasets, metrics, and reference implementations are curated by maintainers and extended via standard contribution workflows. This mirrors established HPC software practices, supporting transparent validation and gradual community convergence on shared quality metrics and tolerances.
@@ 8. Conclusions
We presented astroCAMP as an open benchmark and co-design framework for SKA imaging that couples representative datasets with unified runtime, energy, carbon, and cost metrics. The reported results show: (i) end-to-end limits are dominated by orchestration/synchronization and serial-path effects, not only kernel-level ceilings (Figs. 5, 9, 10); (ii) CPU-only scaling is poor for SKA-sized workloads (Figs. 12, 13); and (iii) sustainability/economic outcomes depend on workload configuration and site context even at equal throughput (Figs. 7, 15). These findings translate into a practical strategy: prioritize pipeline utilization and data movement efficiency, then optimize architecture and scheduling under explicit quality tolerances. We conclude with a community call to: (1) establish science-driven imaging-quality tolerances, (2) expand benchmark datasets, (3) develop reference submissions for current and emerging imaging pipelines, and (4) build an open submission and validation process for the SKA era.